OpenStudy (qwertyzxv):

3. What are the solutions of the system? y = x + 5 y = x^2 - x + 2 A. (3,14) and (-1,0) B. (3,8) and (-1,4) C. (3,8) and (-1,0) D. (-3,2) and (1,6) 4. What are the solutions of the system? y = x^2 - 2x - 8 y = x + 2 A. (5,7) and (-2,0) B. (5,3) and (-2,4) C. (-5,-3) and (2,4) D. (5,3) and (2-4)

jhonyy9 (jhonyy9):

@qwertyzxv you dont like collaborate ?

OpenStudy (mathmale):

y = x + 5 y = x^2 - x + 2 Let's solve this through elimination. Set these 2 equations equal to each other: x+5=x^2-x+2. Subtract (x+5) from both sides and simplify your result. Please share your work here.

OpenStudy (qwertyzxv):

y = x + 5 -x + 5 = y = 10 or 5 .. am i doing this right

OpenStudy (anonymous):

3.) y=x+5 (3,8) 8=3+5 8=8 (-1,4) 4=-1+5 4=4

OpenStudy (mathmale):

First equation: x+5=y Second equation: y=x^2 - x + 2 As before, set the first equation = to the second. y = y. Do the same for x+5 and y^2-x+2: Set them equal to each other. Share your work here, please.

OpenStudy (qwertyzxv):

3 is 3,8 and -1,4 (Thank you so much!! @allyssavillalon) i have no idea what 4. is

jimthompson5910 (jim_thompson5910):

There are a few ways to solve systems like this. One way is to graph each equation on the same xy coordinate grid. Then see where the two graphs cross In this case, the first equation graphs a straight diagonal line. The second equation graphs a parabola See the attached image

jimthompson5910 (jim_thompson5910):

I used desmos to graph https://www.desmos.com/calculator

OpenStudy (anonymous):

4.) Answer is A. y= x^2 - 2x - 8 use (5,7) 7= (5)^2 - 2(5) - 8 7= 25 -10 -8 7= 15 - 8 7=7 y= x+2 (5,7) 7= 5+2 7=7 (-2,0) y=x^2 - 2x - 8 0= (-2)^2 - 2(-2) - 8 0= 4 - (-4) - 8 0= 4+4-(8) 0=8-8 0=0 (-2,0) y=x+2 0= -2 + 2 0=0